OpenStudy (anonymous):

@shifuyanli Marla bought 3 T-shirts for $15.75. How much will it cost her to buy 8 T-shirts at the same rate? $

OpenStudy (anonymous):

okay so what do we do first?

OpenStudy (anonymous):

um..

OpenStudy (anonymous):

15.75*8

OpenStudy (anonymous):

what operation comes first in this problem?

OpenStudy (anonymous):

8-3

OpenStudy (anonymous):

what is 8-3

OpenStudy (anonymous):

$126

OpenStudy (anonymous):

if u dont know it i can ask @TheSmartOne

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@ozzy22 no direct answers

OpenStudy (anonymous):

what now

OpenStudy (agl202):

Do 15.75/3 to get the unit rate then mutliply that by 8

OpenStudy (anonymous):

I do but @ozzy22 gave you the answer so..

OpenStudy (anonymous):

15.75/3=5.25

OpenStudy (anonymous):

yes.

OpenStudy (agl202):

Btw the unit rate was 5.25

OpenStudy (anonymous):

5.25*8=$42

OpenStudy (agl202):

you find the unit rate by dividing 3 from 15.75, 15.75/3=5.25. 5.25 is the cost of 1 shirt and if she got 8 shirts you do 5.25*8 which is 42!

OpenStudy (agl202):

yep thats right!!!

OpenStudy (anonymous):

Solve for x. Express the answer in simplest form.